OpenStudy (anonymous):

What polynomial identity should be used to prove that 162 = (10 + 6)2?

OpenStudy (mathstudent55):

The 2's are exponents, right?

OpenStudy (mathstudent55):

The square of a binomial: \((a + b)^2 = a^2 + 2ab + b^2\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i understand that but i just need to know if i have the right answer

OpenStudy (mathstudent55):

Use a = 10 and b = 6. Substitute those values in for those variables, and evaluate it.

OpenStudy (anonymous):

so would the answer be Square of Binomial

OpenStudy (mathstudent55):

\((a + b)^2 = a^2 + 2ab + b^2 \) \((10 + 6)^2 = 10^2 + 2 \times 10 \times 6 + 6^2\) Now that all you have is numbers, and no more variables, calculate each side.

OpenStudy (anonymous):

is it square of binomial

OpenStudy (mathstudent55):

Yes. If all you need to do is identify the polynomial identity, that's it. If you need to show how it works, then continue to evaluate both sides of the expression above.

OpenStudy (anonymous):

i have anotheer question

OpenStudy (anonymous):

What is the end behavior of the function f(x) = 3x4-x3 + 2x2 + 4x + 5? @mathstudent55

OpenStudy (mathstudent55):

The highest degree term in this polynomial is which term?

OpenStudy (anonymous):

5?

OpenStudy (mathstudent55):

No. The degree of a polynomial is the highest exponent (or sum of exponents) of all terms of the polynomial.

OpenStudy (mathstudent55):

The first term is 3x^4. The degree of the polynomial is 4.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so is it Up on the left, down on the right

OpenStudy (mathstudent55):

Now think of a number raised to the 4th power. When x is negative, and getting more negative, as it approaches negative infinity, what happens to x^4?

OpenStudy (anonymous):

so am i right

OpenStudy (mathstudent55):

Not so fast. We haven't looked at when x is positive and approaches positive infinity.

OpenStudy (mathstudent55):

Once again, When x is negative, and getting more negative, as it approaches negative infinity, what happens to x^4?

OpenStudy (anonymous):

it goes down

OpenStudy (anonymous):

so am i right

OpenStudy (mathstudent55):

If you take a negative number and raise it to the 4th power, what do you get?

OpenStudy (anonymous):

Up on the left, up on the right Up on the left, down on the right Down on the left, up on the right Down on the left, down on the right

OpenStudy (anonymous):

i have to go soon so can you tell me which ones right

OpenStudy (mathstudent55):

As the x gets closer to negative infinity, what happens to \(x^4\)? \((-1)^4 = 1\) \((-2)^4 = 16\) \((-3)^4 = 81\) \((-4)^4 = 256\) \((-10)^4 = 10,000\)

OpenStudy (anonymous):

please i need to go soon

OpenStudy (mathstudent55):

It gets larger. The same thing happens when x gets closer to positive infinity. x^4 gets larger. The graph of this polynomial will look like this: